A buffer solution is 0.20M in acetic acid and in sodium acetate. Calculate the change in pH up on adding 1.0ml of 0.10M HCl acid to 10 ml of this solution. ?
1
Expert's answer
2020-09-24T13:50:14-0400

The ICE approach for the equilibrium yields;

"CHCO_2H+H_2O \leftrightharpoons H_3O^{+}+CH_3CO_2^-"

The initial concentration of "CH_3CO_2H" "=0.20"

Change "=x"

Equilibrium concentration"= 0.20-x"

The initial concentration of "CH_3CO_2^-"

."=0.20"

Change"=x"

Equilibrium concentration"= 0.20+x"

Substituting the equilibrium concentration terms into the "K_a" expression, assuming "x\lll" "0.20" and solving the simplified equation for "x" yields

"x= 1.8\times 10^{-5}"

"[H_3O^+]=0+x=1.8\times 10^{-5}" "pH=-log [H_3O^+]=-log (1.8\times 10^{-5}M)=4.74"